OpenStudy (anonymous):

What is the slope of the line that passes through the pair of points? (1, 7), (10, 1) A. 3/2 B. -3/2 C. -2/3 D. 2/3

OpenStudy (anonymous):

\[m = (y _{2}-y _{1})/(x _{2}-x_{1})\]

OpenStudy (texaschic101):

it can be solved by using the slope formula as suggested by ASNinja. slope(m) = (y2 - y1) / (x2 - x1) (1,7) x1 = 1 and y1 = 7 (10,1) x2 = 10 and y2 = 1 now just sub slope(m) = (1 - 7) / (10 - 1) --- can you finish ??

OpenStudy (anonymous):

y-y1/x-x1 = y2-y1/x2-x1 y-7/x-1 = 1-7/10-1 y-7/x-1 = -6/9 y-7 = -6x/9 +6/9 y=-6x/9 + 23/3 i think im doing that right if not let me know please

OpenStudy (texaschic101):

what are you doing ?? slope(m) = (y2 - y1) / (x2 - x1) and when we subbed our points in, we got : slope(m) = (1 - 7) / (10 - 1) slope(m) = -6/9 which reduces to - 2/3

OpenStudy (anonymous):

oh right i see what u did.. sorry I get confused with math